# AdaLRS: Loss-Guided Adaptive Learning Rate Search for Efficient Foundation Model Pretraining

### v2

@@ Abstract
Learning rate is widely regarded as crucial for effective foundation model pretraining. Recent research explores and demonstrates the transferability of learning rate configurations across varying model and dataset sizes, etc. Nevertheless, these approaches are constrained to specific training scenarios and typically necessitate extensive hyperparameter tuning on proxy models. In this work, we propose AdaLRS, a plug-in-and-play adaptive learning rate search algorithm that conducts online optimal learning rate search via optimizing loss descent velocities. We provide experiment results to show that the optimization of training loss and loss descent velocity in foundation model pretraining are both convex and share the same optimal learning rate. Relying solely on training loss dynamics, AdaLRS involves few extra computations to guide the search process, and its convergence is guaranteed via theoretical analysis. Experiments on both LLM and VLM pretraining show that AdaLRS adjusts suboptimal learning rates to the neighborhood of optimum with marked efficiency and effectiveness, with model performance improved accordingly. We also show the robust generalizability of AdaLRS across varying training scenarios, such as different model sizes, training paradigms, and base learning rate scheduler choices.
* Project lead†Email corresponding
@@ 1 Introduction
  • We provide experimental analysis to demonstrate that the optimization of training loss and loss slopes w.r.t. LR are convex for foundation model pretraining, and they share the same optimum.
@@ 2.1 Formulations of AdaLRS Algorithm
$\eta_{t+k}=\begin{cases}\alpha^{\prime}\eta_{t}&\text{if }v(\alpha^{\prime}\eta_{t})>v(\eta_{t})+2e\quad(\text{loss slope {increases} $\uparrow$}),\\ \beta^{\prime}\eta_{t}&\text{if }v(\beta^{\prime}\eta_{t})<v(\eta_{t})-2e\quad(\text{loss slope {decreases} $\downarrow$}),\\ \eta_{t}&\text{otherwise.}\end{cases}$ (1)
$v(\cdot)$ indicates the estimated loss curve slope obtained from a $k$-step window, while $e$ stands for the estimation error between $v$ and the true loss descent velocity $V$. $\alpha^{\prime}$ and $\beta^{\prime}$ are rectified LR scaling factors which satisfy $\alpha^{\prime}=\max(\lambda^{t}\alpha,1),\beta^{\prime}=\frac{1}{\max(\lambda^{t}\beta,1)}$, where $\alpha,\beta>1$ are co-prime integers and $\lambda\in(0,1)$ is a decay factor. We validate the co-prime design of LR scaling factors in Appendix B.
@@ 2.2.1 Theoretical Hypotheses
This is a relatively strong assumption, and therefore we conduct a series of pilot studies to support the it. Specifically, we conduct small-scale pretraining experiments in LLM and VLM pretraining. For LLM pretraining, we use Qwen2.5-1.5B and Qwen2.5-7B qwen2.5 models with randomly initialized parameters, and train them from scratch on 8M samples from the SlimPajama cerebras2023slimpajama dataset. A series of exponentially larger learning rates in $[2e^{-6},8e^{-3}]$ are applied to investigate the dynamics of training loss and loss curve slope under varying LR settings. For VLM pretraining, we conduct experiments on a 2B model with the architecture described in SAIL-VL dong2025scalable . Approximately 8M data consisting of caption and OCR samples are used for model training, with learning rates configured in $[2e^{-5},2]$.
@@ 2.2.2 Convergence
In practice, the value of $\lambda$ is set relatively large, ensuring that the learning rate $\eta$ converges to the neighborhood of $\eta^{*}$ at the current LR adjustment magnitude before the scaling factor becomes too small. Since the desired LR adjustment is guaranteed by Proposition 2.2, we begin with $\frac{\eta^{*}-e}{\alpha^{\prime}}<\eta_{t}<\frac{\eta^{*}+e}{\beta^{\prime}}$. In this way, the learning rate $\eta_{t}$ either oscillates within this range, or attempts to search the learning rate in $[\frac{\eta^{*}+e}{\beta^{\prime}},\frac{\eta^{*}+e}{\beta^{\prime}}\alpha^{\prime}]$ and $[\frac{\eta^{*}-e}{\alpha^{\prime}}\beta^{\prime},\frac{\eta^{*}-e}{\alpha^{\prime}}]$ respectively for higher or lower learning rate values. Since we have $\frac{\eta^{*}+e}{\beta^{\prime}}>\eta^{*}+e$ always holds true, the LR upscaling attempts will trigger LR downscaling according to Proposition 2.2, maintaining the LR value given in Proposition 2.3. Similarly, learning rate downscaling attempts will be rejected by Proposition 2.2.
@@ 3.1 Experiment Setup
To demonstrate the effectiveness of the AdaLRS algorithm, we set different learning rates for each training task, i.e., learning rates too small, too large, and appropriate for pretraining. For each setting, we train a baseline model and a model optimized by AdaLRS for fair comparison. We set the upscaling factor $\alpha$, downscaling factor $\beta$, and decaying factor $\lambda$ as $3$, $2$, and $0.99$ in all experiments for LR adjustment effectiveness. Approximately 120B and 160B tokens are used for LLM and VLM pretraining, with roughly 10,000 and 20,000 910B NPU hours consumed for 2B and 7B model pretraining experiments. A cosine learning rate scheduler loshchilov2016sgdr is applied in the main experiments. Detailed model training recipes are elaborated in Table 1.
@@
Evaluation. To demonstrate the effectiveness of the proposed AdaLRS algorithm, we show the learning rate search and training loss dynamics of our foundation model training experiments. For LLM pretraining, we quantify the performance advantage of models trained with AdaLRS with final training loss and perplexity (PPL) computed on SlimPajama cerebras2023slimpajama train, validation, and test splits. For VLM experiments, however, there is no widely accepted benchmark or metric for pretrained VLM evaluation. Therefore, we tune the pretrained VLM on 3M samples from Infinity-MM gu2024infinity stage3 data with all parameters unfrozen, and use a series of publicly available VLM benchmarks for model evaluation, such as LLaVABench liu2024visual , MMVet yu2024mm , MMStar chen2024we , DocVQA mathew2021docvqa , OCRBench liu2024ocrbench , TextVQA singh2019towards , DetailCaps-4870 dong2024benchmarking , etc.
@@ 3.2 Main Results
Training PPL $13.72/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}12.65}$ $183.28/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}205.00}$ $12.88/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}33.21}$ $10.84/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}10.88}$ $158.54/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}165.67}$ $10.76/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}15.49}$
@@
Fit LR $39.5/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}38.5}$ $34.58/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}32.02}$ $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}48.67/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}49.53}$ $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}77.99/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}78.00}$ $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}718/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}735}$ $64.89/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}63.74}$ $55.68/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}55.30}$ $52.24/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}51.82}$
@@
Large LR $36.8/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}35.7}$ $44.47/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}44.33}$ $31.47/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}30.50}$ $57.53/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}57.42}$ $631/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}606}$ $60.32/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}58.08}$ $49.02/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}47.08}$ $49.44/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}42.96}$
@@
Small LR $44.3/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}39.2}$ $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}48.67/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}49.20}$ $36.15/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}30.23}$ $77.75/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}77.47}$ $730/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}689}$ $64.85/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}57.86}$ $56.65/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}53.51}$ $51.81/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}51.70}$
@@
AdaLRS Enhances Pretrained Foundation Model Performance Markedly. We provide PPL evaluation results for the pretrained LLMs in Table 2 to demonstrate the model performance advantage of models trained with AdaLRS. AdaLRS not only surpasses baseline LLMs pretrained with suboptimal learning rates, but also achieves comparable training loss and evaluation PPL compared with Fit LR baselines. It is worth noticing that AdaLRS improves model performance significantly for excessively small LRs, even surpassing the Fit LR baselines in 7B LLM pretraining. This observation implies that starting with a relatively small LR setting, AdaLRS can potentially optimize both the learning rate and model performance for unexplored pretraining tasks within a single run.

### v3

@@ Abstract
Learning rate is widely regarded as crucial for effective foundation model pretraining. Recent research explores and demonstrates the transferability of learning rate configurations across varying model and dataset sizes, etc. Nevertheless, these approaches are constrained to specific training scenarios and typically necessitate extensive hyperparameter tuning on proxy models. In this work, we propose AdaLRS, a plug-in-and-play adaptive learning rate search algorithm that conducts online optimal learning rate search via optimizing loss descent velocities. We provide theoretical and experimental analyzes to show that foundation model pretraining loss and its descent velocity are both convex and share the same optimal learning rate. Relying solely on training loss dynamics, AdaLRS involves few extra computations to guide the search process, and its convergence is guaranteed via theoretical analysis. Experiments on both LLM and VLM pretraining show that AdaLRS adjusts suboptimal learning rates to the neighborhood of optimum with marked efficiency and effectiveness, with model performance improved accordingly. We also show the robust generalizability of AdaLRS across varying training scenarios, such as different model sizes, training paradigms, base learning rate scheduler choices, and hyperparameter settings.
†Email corresponding
@@ 1 Introduction
  • We provide theoretical and experimental analysis to demonstrate that the foundation model pretraining loss and its slop w.r.t. LR are convex and share the same optimum.
@@ 2.1 Formulations of AdaLRS Algorithm
[…]
$v(\cdot)$ indicates the estimated loss curve slope obtained from a $k$-step window, while $e$ stands for the estimation error between $v$ and the true loss descent velocity $V$. $\alpha^{\prime}$ and $\beta^{\prime}$ are rectified LR scaling factors which satisfy $\alpha^{\prime}=\max(\lambda^{t}\alpha,1),\beta^{\prime}=\frac{1}{\max(\lambda^{t}\beta,1)}$, where $\alpha,\beta>1$ are two multiplicatively independent real numbers and $\lambda\in(0,1)$ is a decay factor. We validate the multiplicatively independent design of LR scaling factors (for all integers $m,n$, $\alpha^{m}=\beta^{n}\implies m=n=0$) in Appendix B.
@@ 2.2.1 Theoretical Hypotheses
This is a relatively strong assumption, and therefore we conduct both theoretical and experimental analyses to support it. Consider a simplified foundation model pretraining task with a constant learning rate and SGD optimizer, we formulate the update rule for model parameter $\psi$ as:
$\psi_{t+1}=\psi_{t}-\eta\nabla L_{t},$ (3)
where $L_{t}$ is the training loss at time step $t$. The expected loss descent velocity per step (using Taylor expansion) is:
$\mathbb{E}\left[L_{t+1}-L_{t}\right]\approx-\eta\|\nabla L_{t}\|^{2}+\frac{C_{Lip}}{2}\eta^{2}\|\nabla L_{t}\|^{2},$ (4)
where $C_{Lip}$ is the Lipschitz constant of the gradient. When the learning rate is small, the first term dominates the expectation, and a smaller $\eta$ leads to a smaller decrease in loss. When the learning rate is large, on the other hand, the second term cannot be neglected. It contributes to suppress the expected loss descent value. Differentiating w.r.t. $\eta$ and setting to zero for the extreme:
$\frac{\partial}{\partial\eta}\left(-\eta+\frac{C_{Lip}}{2}\eta^{2}\right)=0\implies\eta^{*}=\frac{1}{C_{Lip}}.$ (5)
This shows that the loss descent velocity w.r.t. $\eta$ is a convex function. It is intuitive to infer that a higher loss descent velocity leads to a lower training loss, and vice versa, which proves that the optimization of training loss is also a convex problem.
Apart from theoretical analyses, we also conduct small-scale pretraining experiments in LLM and VLM pretraining to verify this hypothesis. For LLM pretraining, we use Qwen2.5-1.5B and Qwen2.5-7B qwen2.5 models with randomly initialized parameters, and train them from scratch on 8M samples from the SlimPajama cerebras2023slimpajama dataset. A series of exponentially larger learning rates in $2e^{-6},8e^{-3}]$ are applied to investigate the dynamics of training loss and loss curve slope under varying LR settings. For VLM pretraining, we conduct experiments on a 2B model with the architecture described in SAIL-VL dong2025scalable. Approximately 8M data consisting of caption and OCR samples are used for model training, with learning rates configured in $2e^{-5},2]$.
@@ 2.2.2 Convergence
In practice, the value of $\lambda$ is set relatively large, ensuring that the learning rate $\eta$ converges to the neighborhood of $\eta^{*}$ at the current LR adjustment magnitude before the scaling factor becomes too small. Since the desired LR adjustment is guaranteed by Proposition 2.2, we begin with $\frac{\eta^{*}-e}{\alpha^{\prime}}<\eta_{t}<\frac{\eta^{*}+e}{\beta^{\prime}}$. In this way, the learning rate $\eta_{t}$ either oscillates within this range, or attempts to search the learning rate in $\frac{\eta^{*}+e}{\beta^{\prime}},\frac{\eta^{*}+e}{\beta^{\prime}}\alpha^{\prime}]$ and $\frac{\eta^{*}-e}{\alpha^{\prime}}\beta^{\prime},\frac{\eta^{*}-e}{\alpha^{\prime}}]$ respectively for higher or lower learning rate values. Since we have $\frac{\eta^{*}+e}{\beta^{\prime}}>\eta^{*}+e$ always holds true, the LR upscaling attempts will trigger LR downscaling according to Proposition 2.2, maintaining the LR value given in Proposition 2.3. Similarly, learning rate downscaling attempts will be rejected by Proposition 2.2.
@@ 3.1 Experiment Setup
To demonstrate the effectiveness of the AdaLRS algorithm, we set different learning rates for each training task, i.e., learning rates too small, too large, and appropriate for pretraining. For each setting, we train a baseline model and a model optimized by AdaLRS for fair comparison. We set the upscaling factor $\alpha$, downscaling factor $\beta$, and decaying factor $\lambda$ as $3$, 22, and $0.99$ in all experiments for LR adjustment effectiveness. Approximately 120B and 160B tokens are used for LLM and VLM pretraining, with roughly 10,000 and 20,000 910B NPU hours consumed for 2B and 7B model pretraining experiments. A cosine learning rate scheduler loshchilov2016sgdr is applied in the main experiments. Detailed model training recipes are elaborated in Table 1.
@@
Evaluation. To demonstrate the effectiveness of the proposed AdaLRS algorithm, we show the learning rate search and training loss dynamics of our foundation model training experiments. For LLM pretraining, we quantify the performance advantage of models trained with AdaLRS with final training loss and perplexity (PPL) computed on SlimPajama cerebras2023slimpajama train, validation, and test splits. We also conduct a lightweight SFT on 6M samples from the Infinity-Instruct dataset li2025infinity, and evaluate the downstream model performance on open-ended generation tasks such as Alpaca-Gen and KNIGHT-Gen yang2025improving. For VLM experiments, however, there is no widely accepted benchmark or metric for pretrained VLM evaluation. Therefore, we tune the pretrained VLM on 3M samples from Infinity-MM gu2024infinity stage3 data with all parameters unfrozen, and use a series of publicly available VLM benchmarks for model evaluation, such as LLaVABench liu2024visual, MMVet yu2024mm, MMStar chen2024we, DocVQA mathew2021docvqa, OCRBench liu2024ocrbench, TextVQA singh2019towards, DetailCaps-4870 dong2024benchmarking, etc.
@@ 3.2 Main Results
[…]
@@
Alpaca-Gen ${\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}17.29}/18.56$ $7.11/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}6.09}$ $17.10/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}15.40}$ $21.76/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}21.61}$ $6.35/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}5.55}$ $21.15/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}20.68}$
KNIGHT-Gen ${\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}10.13}/11.29$ $3.96/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}2.02}$ $10.81/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}8.36}$ $13.62/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}13.35}$ $4.08/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}3.82}$ $13.53/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}12.29}$
@@
Fit LR $39.5/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}38.5}$ $34.58/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}32.02}$ $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}48.67/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}49.53}$ $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}77.99/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}78.00}$ $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}718/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}735}$ $64.89/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}63.74}$ $55.68/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}55.30}$ $56.16/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}55.80}$
@@
Large LR $36.8/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}35.7}$ $31.47/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}30.50}$ $44.47/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}44.33}$ $57.53/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}57.42}$ $631/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}606}$ $60.32/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}58.08}$ $49.02/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}47.08}$ $48.96/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}47.67}$
@@
Small LR $44.3/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}39.2}$ $36.15/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}30.23}$ $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}48.67/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}49.20}$ $77.75/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}77.47}$ $730/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}689}$ $64.85/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}57.86}$ $56.65/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}53.51}$ $57.34/{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}53.77}$
@@
AdaLRS Enhances Pretrained Foundation Model Performance Markedly. To demonstrate the model performance advantage of training with AdaLRS, we provide PPL and downstream task evaluation results for LLM experiments in Table 2. AdaLRS not only surpasses baseline LLMs pretrained with suboptimal learning rates, but also achieves comparable model performance with Fit LR baselines. It is worth noticing that AdaLRS improves model performance significantly for excessively small LRs, even surpassing the Fit LR baselines in 7B LLM pretraining. This observation implies that starting with a relatively small LR setting, AdaLRS can potentially achieve both the optimal learning rate and model performance for unexplored pretraining tasks within a single run.